Corporate bonds issued by Johnson Corporation currently yield 11.5%. Municipal bonds of equal risk currently yield 6%. At what tax rate would an investor be indifferent between these two bonds? Round your answer to two decimal places.
Answer #1

corporate bond yield*(1-tax rate)=municipal bond yield

11.5*(1-tax rate)=6

1-tax rate=(6/11.5)

tax rate=1-(6/11.5)

which is equal to

=47.83%(Approx).